Find 3 consecutive positive odd integers such that the product of the first and second integers is eight more than the third integer.
Answer by TimothyLamb(4379)

z = y + 2
y = x + 2
xy = z + 8
---
z = (x + 2) + 2
z = x + 4
--
xy = z + 8
x(x + 2) = z + 8
x(x + 2) = (x + 4) + 8
xx + 2x = x + 12
xx + x - 12 = 0
---
the above quadratic equation is in standard form, with a=1, b=1 and c=-12
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 1 -12
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 3
x = -4
---
the negative root doesn't fit the problem statement, so use the positive root:
---
answer:
x = 3
y = 5
z = 7
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php